### v1

# Best Ergodic Averages via Optimal Graph Filters in Reversible Markov Chains

Abstract
In this paper, we address the problem of finding the best ergodic or Birkhoff averages in the ergodic theorem to ensure rapid convergence to a desired value, using graph filters. Our approach begins by representing a function on the state space as a graph signal, where the (directed) graph is formed by the transition probabilities of a reversible Markov chain. We introduce a concept of graph variation, enabling the definition of the graph Fourier transform for graph signals on this directed graph. Viewing the iteration in the ergodic theorem as a graph filter, we recognize its non-optimality and propose three optimization problems aimed at determining optimal graph filters. These optimization problems yield the Bernstein, Chebyshev, and Legendre filters. Numerical testing reveals that while the Bernstein filter performs slightly better than the traditional ergodic average, the Chebyshev and Legendre filters significantly outperform the ergodic average, demonstrating rapid convergence to the desired value.
Index Terms:
Ergodic theorem, graph filters, Bernstein-Chebyshev-Legendre polynomials.
I Introduction
The ergodic theorem is a cornerstone in the study of dynamical systems and stochastic processes, providing a foundation for understanding long-term average behavior. Traditional ergodic or Birkhoff averages are essential tools in this significant result, but their convergence rates can be slow. This paper seeks to enhance these averages using graph signal processing techniques, particularly through the application of graph filters.
Our approach begins by conceptualizing a function on the state space as a graph signal, where the (directed) graph, on which a graph signal can be defined, is constructed based on the transition probabilities of a reversible Markov chain. By introducing a notion of graph variation, we can define the graph Fourier transform for graph signals on this directed graph. This new perspective allows us to interpret the iteration in the ergodic theorem as a graph filter. However, this standard ergodic graph filter is not optimal in terms of convergence speed. To address this issue, we formulate three distinct optimization problems, each with a different objective, to derive the optimal graph filters. The solutions to these problems result in the Bernstein, Chebyshev, and Legendre polynomial filters.
We then validate our theoretical findings through numerical experiments. The results indicate that the Bernstein filter offers a slight improvement over the ergodic average. More notably, the Chebyshev and Legendre filters exhibit significantly better performance, achieving rapid convergence to the desired value. These findings highlight the effectiveness of our proposed graph signal processing approach in optimizing ergodic averages.
Our paper intersects with two main lines of research: faster algorithm design via approximation theory and optimal Birkhoff averages in ergodic theorem for deterministic dynamical systems. The first domain aims to accelerate algorithms by approximating some well-known functions, which appear in the computation of outputs of these algorithms, using polynomials [1]. The second focuses on selecting optimal ergodic averages to enhance the convergence rates in ergodic theorem for deterministic dynamical systems [2].
In algorithm design, approximation theory plays a crucial role in developing efficient computational methods. Early applications include the conjugate gradient method [3, 4] and the Lanczos method [5, 6], both of which are used to solve systems of linear equations. These techniques significantly reduce the number of computations compared to traditional methods like Gaussian elimination. Furthermore, approximation theory also extends its usage to complexity theory. For instance, Chebyshev polynomials have been used to enhance quantum search algorithms [7]. Particularly relevant to our study is the relatively recent monograph [1], which explores how approximating well-known functions with polynomials or rational functions can accelerate algorithms. This monograph studies techniques for approximating functions such as $x^{s}$, $x^{-1}$, and $e^{-x}$ and investigates how these results play an important role in the design of fast algorithms for some well-known problems. In particular, these techniques facilitate more efficient computation of fundamental objects like $A^{s}v$, $A^{-1}v$, $\exp(-A)v$, eigenvalues, and eigenvectors, which are integral to many spectral algorithms. In fact, numerous algorithms can be reduced to the calculation of these objects, which have been shown to significantly expedite essential computations such as random walk simulations, graph partitioning, and solving systems of linear equations.
In this monograph, the primary objective is to approximate the aforementioned functions with low-degree polynomials given a specified maximum error, without an optimization component. In contrast, our approximations are performed optimally by formulating three optimization problems with distinct objectives to identify the best polynomial, resulting in three well-known polynomials in approximation theory. This approach allows practitioners to select the most suitable polynomial based on the specific application by considering the appropriate objective. Furthermore, while the monograph deals with fixed functions, our approach specifies the function to approximate only over a sub-interval of its domain and on a single isolated point. This distinction makes our problem significantly different from those addressed in the monograph. Additionally, our paper introduces a novel perspective to these approximation problems by interpreting them as graph filters. This enables us to use established signal processing methods, a viewpoint that is absent in the monograph.
In the context of ergodic theory for deterministic dynamical systems, the Birkhoff ergodic theorem states that, for ergodic systems, the time averages of a function along a trajectory of length $T$ will converge to the space average as $T$ approaches infinity. However, this convergence is often slow. To address this, researchers have explored using non-uniform weights for averages, instead of uniform averages that equally weight each point along the trajectory. This method is particularly relevant to our problem formulation. In [2], the authors address this issue by assigning significantly lower weights to the initial and final points of the trajectory compared to those near the midpoint using a bump function. They demonstrate that, in quasiperiodic dynamical systems, these weighted averages achieve much faster convergence, provided that the function $f$ has sufficient differentiability. The studies in [8, 9, 10] also discuss the problem of accelerating the convergence of weighted ergodic sums, yet they do not specify the exact rates of convergence as presented in [2]. In [11], for a certain choice of weights, a convergence rate is obtained which is worse than the one in [2]. In [12], authors established the polynomial and exponential convergence rate of weighted Birkhoff averages of irrational rotations on tori, which improves the result in [2]. In these works, there is no optimization component in the problem formulation as opposed to our case. Namely, it is established that if the function generating the weights satisfies certain properties, the weighted Birkhoff averages converge to the space average at some speed. However, they do not claim that this is the optimal convergence rate achievable, nor do they assert that the weights are optimal.
In this direction, the most pertinent work related to our study is [13]. In this work, the authors expand the function $f$, whose integral is to be computed, using basis functions and view the weighted Birkhoff averages as a filter on the coefficients of this expansion. They observe that a filter tuned to specific frequencies can be significantly more effective than an arbitrary bump function, as considered in [2]. This approach aligns with the method we take in this paper. However, there are notable distinctions. First, [13] considers deterministic dynamical systems on abstract topological spaces, whereas we focus on stochastic dynamics given by a Markov chain on a finite set. Additionally, in [13], there are infinitely many frequency components, and so, the function approximated via filters differs from ours. Furthermore, the basis functions corresponding to the frequencies are not related to the eigenfunctions of the dynamical map, meaning there is no operator-theoretic perspective or spectral analysis of the dynamical map. Finally, there is no underlying graph structure in their work, unlike in our case. It is worth mentioning that our work was developed independently of [13]. Consequently, our motivation for employing a filtering approach to achieve fast convergence in the ergodic theorem for Markov chains stems entirely from graph signal processing theory. Nevertheless, it is fascinating to see that such a natural approach is also being embraced by the dynamical systems community.
We note that even if one views optimal Birkhoff averages in deterministic dynamical systems via an operator theory perspective, a complication still arises as the operators are generally infinite-dimensional, and so their spectra are not well-behaved (isolated) as in the finite-dimensional case. Hence, graph signal processing techniques cannot be applied. Indeed, an important research problem in this direction is as follows: consider a Markov chain with an abstract state space. Then the transition probability can be viewed as an infinite-dimensional operator, where the spectrum may contain infinitely many elements. In this case, the eigenvalue $1$ is probably not isolated, and so we cannot directly apply graph filters as we cannot easily attenuate eigenvalues that are closer to $1$. In this case, we can first approximate the operator with a finite-dimensional one and then apply our graph filters. This is indeed a research direction we want to pursue in the future.
This paper is structured as follows: Section II reviews reversible Markov chains. Section III details our perspective about the ergodic theorem, including the construction of the graph signal framework and the definition of graph Fourier transform. Sections IV, V, and VI present the formulations and solutions of the optimization problems that lead to Bernstein, Chebyshev, and Legendre polynomial filters, respectively. Section VII discusses our numerical experiments and results. Finally, Section VIII offers concluding remarks.
II Reversible Markov Chains
Let ${\mathsf{X}}$ be a finite set and let $P:{\mathsf{X}}\rightarrow{\mathcal{P}}({\mathsf{X}})$ be a transition probability of a Markov chain on ${\mathsf{X}}$. By an abuse of notation, we also denote the transition matrix via $P$, and so, $P\in\mathbb{R}^{{\mathsf{X}}\times{\mathsf{X}}}$. Therefore, $P(x,y)=\mathop{\rm Pr}\{X_{t+1}=y|X_{t}=x\}$ for all $x,y\in{\mathsf{X}}$. We assume that $P$ is irreducible. In this case, it is known that there exists a unique stationary distribution $\pi$ of $P$ (see [14, Corollary 1.17]) and $\pi(x)>0$ for all $x\in{\mathsf{X}}$ (see [14, Proposition 1.19]).
Definition 1.
The transition probability $P$ is said to be reversible if there exists a probability distribution $\pi$ on ${\mathsf{X}}$ such that
$\pi(x)P(x,y)=\pi(y)P(y,x)\,\,\,\,\forall x,y\in{\mathsf{X}}.$
The above equation is called detailed balance equation.
By [14, Proposition 1.20], any probability distribution that satisfies the detailed balance equation is a stationary distribution of $P$:
$\big{[}\pi(x)P(x,y)=\pi(y)P(y,x)\,\,\,\,\forall x,y\in{\mathsf{X}}\big{]}\implies\big{[}\pi P=\pi\big{]}.$
Moreover, the corresponding Markov chain $\{X_{t}\}$ with initial distribution $\pi$ satisfies the following: for any $n$, we have
$\mathop{\rm Pr}\{X_{0}=x_{0},X_{1}=x_{1},\ldots,X_{n}=x_{n}\}$
$=\pi(x_{0})P(x_{0},x_{1})\ldots P(x_{n-1},x_{n})$
$=\pi(x_{n})P(x_{n},x_{n-1})\ldots P(x_{1},x_{0})$
$=\mathop{\rm Pr}\{X_{0}=x_{n},X_{1}=x_{n-1},\ldots,X_{n}=x_{0}\}.$
Hence, the time reversed version of the Markov chain $\{X_{t}\}$ is indistinguishable from the original Markov chain.
Example 1.
Random Walk on Graphs
Let ${\mathcal{G}}=({\mathsf{X}},E)$ be an undirected, simple, and connected graph with the vertex set ${\mathsf{X}}$ and the edge set $E$. A random walk on ${\mathcal{G}}$ is described as follows: if $X_{t}=x$ is the current position, then the next position is chosen uniformly from the set of neighbors of $x$. Hence, if $d(x)$ is the degree of $x$, then
$P(x,y)=\begin{cases}\frac{1}{d(x)}&\text{if}\,\,\{x,y\}\in E\\ 0&\,\,\text{otherwise}\end{cases}$
Since the graph is connected, the corresponding Markov chain is irreducible. It is also straightforward to see that the Markov chain is reversible with respect to
$\pi(x)=\frac{d(x)}{2|E|},\,\,x\in{\mathsf{X}},$
where $|E|$ is the number of edges.
Example 2.
Glauber Chains
In this example, $V$ and $S$ are finite sets. We can think of $V$ as the set of vertices of a graph and $S$ is the set of possible values that can be taken by these vertices. Our state space ${\mathsf{X}}$ is a subset of $V^{S}$. Let $\pi$ be a probability distribution on ${\mathsf{X}}$. In statistical physics, $\pi$ is called a Gibbs distribution.
The Glauber dynamics for $\pi$ is a reversible Markov chain with state space ${\mathsf{X}}$ and stationary distribution $\pi$. The transition probability is given as follows: if $X_{t}=x$ is the current state, then we generate a point (vertex) $v$ from $V$ uniformly and pick the next state randomly via
$X_{t+1}=y\sim\pi(y|{\mathsf{X}}(x,v))=\begin{cases}\frac{\pi(y)}{\pi({\mathsf{X}}(x,v))}&\text{if}\,\,y\in{\mathsf{X}}(x,v)\\ 0&\text{otherwise}.\end{cases}$
Here, ${\mathsf{X}}(x,v)\coloneqq\{y\in{\mathsf{X}}:y(w)=x(w)\,\,\forall\,w\neq v\}$. Verbally, a new state is chosen according to the probability distribution $\pi$ conditioned on the set of states equal to the current state $x$ at all vertices different from $v$.
Define ${\mathsf{X}}(x)\coloneqq\bigcup_{v\in V}{\mathsf{X}}(x,v)$. Then, the transition probability can be defined as follows:
$P(x,y)=\begin{cases}\frac{1}{|V|}\pi(y|{\mathsf{X}}(x,v))&\text{if}\,\,y\in{\mathsf{X}}(x)\,\text{and}\,y\neq x\\ \sum_{v\in V}\frac{1}{|V|}\pi(y|{\mathsf{X}}(x,v))&\text{if}\,\,y=x\\ 0&\text{otherwise}.\end{cases}$
Note that if $\pi(x)>0$ for all $x\in{\mathsf{X}}$, then $P$ is irreducible. Moreover, one can also prove that $P$ and $\pi$ satisfy the detailed balance condition. Hence, $\pi$ is a stationary distribution of $P$ and $P$ is reversible.
A specific Glauber dynamics can be constructed if $\pi$ is a Gibbs distribution of the following form:
$\pi(x)=\frac{e^{-\beta H(x)}}{Z(\beta)},$
where $H(x)$ is the energy of the state $x$ and $Z(\beta)\coloneqq\sum_{x\in{\mathsf{X}}}e^{-\beta H(x)}$ is the partition function. For instance, if $S=\{-1,+1\}$ and $V$ is the vertex set of some finite graph, then
$H(x)=-\sum_{u,w\in V:u\sim w}J_{v,w}\,x(v)x(w),$
where $J_{v,w}>0$. Then, the transition probability is given by
[…]
where $S(x,w)\coloneqq\sum_{u:u\sim w}J_{u,w}\,x(u)$.
II-A Eigenvalues of Reversible Markov Chains
Let $P$ be a reversible Markov chain with respect to $\pi$. Then, we define the following inner product on $\mathbb{R}^{{\mathsf{X}}}$:
$\langle f,g\rangle_{\pi}\coloneqq\sum_{x\in{\mathsf{X}}}f(x)g(x)\pi(x).$
The following result follows from [14, Lemmas 12.1 and 12.2].
Lemma 1.
Let $P$ be a reversible Markov chain with respect to $\pi$. Then, we have
  • (1)
    $P$ has real eigenvalues. We denote this set via $\sigma(P)$.
  • (2)
    $\sigma(P)\subset[-1,1]$.
  • (3)
    There exists orthonormal basis of real-valued eigenfunctions $\{f_{x}\}_{x\in{\mathsf{X}}}$ corresponding to real eigenvalues $\{\lambda_{x}\}_{x\in{\mathsf{X}}}$ for inner product space $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle_{\pi})$.
The spectral gap of a reversible Markov chain $P$ is defined as $\gamma\coloneqq 1-\lambda_{2}$, where $\lambda_{2}$ is the second largest eigenvalue of $P$. The spectral gap, or a lower bound to it, is important in the optimal graph filter design. We close this subsection by stating the Birkhoff ergodic theorem for irreducible Markov chains.
Theorem 1.
Birkhoff Ergodic Theorem
Let $P$ be an irreducible Markov chain with a unique stationary distribution $\pi$. Then, for any real-valued function $f$ on ${\mathsf{X}}$, we have
$\lim_{t\rightarrow\infty}\mathbb{E}\left[\frac{1}{t}\sum_{k=0}^{t-1}f(X_{t})\bigg{|}X_{0}=x\right]=\sum_{z\in{\mathsf{X}}}f(z)\pi(z)\,\,\forall x\in{\mathsf{X}},$
where $\{X_{t}\}$ is Markov chain with transition probability $P$.
Note that we can write
$\mathbb{E}\left[\frac{1}{t}\sum_{k=0}^{t-1}f(X_{t})\bigg{|}X_{0}=x\right]=\frac{1}{t}(\mathbb{Id}+P+\ldots+P^{t-1})f(x).$
Hence, ergodic theorem implies that $\frac{1}{t}(\mathbb{Id}+P+\ldots+P^{t-1})f$ converges to the constant vector $\left(\sum_{z\in{\mathsf{X}}}f(z)\pi(z)\right){\bf 1}$ pointwise. Let $\pi(f)=\sum_{z\in{\mathsf{X}}}f(z)\pi(z)$. The iteration in ergodic theorem
$\frac{(\mathbb{Id}+P+\ldots+P^{t-1})}{t}f$
can be considered as a low-pass graph filter for the graph signal $f$, as we will elaborate on in the next section. Therefore, in the remainder of this paper, we aim to develop algorithms to identify the optimal low-pass graph filter that achieves convergence to $\pi(f){\bf 1}$ faster than the aforementioned iteration.
III Graph Signal Processing
In this section, ${\mathcal{G}}=({\mathsf{X}},E)$ denotes a simple (directed) graph with a finite vertex set ${\mathsf{X}}$ and an edge set $E$. Therefore, if $(x,y)\in E$, there is a directed edge from $x$ to $y$, denoted as $x\rightarrow y$. Each directed edge in this graph has an associated weight, denoted by $P(x,y)\in\mathbb{R}_{+}$. Since we are interested in Markov chains, we assume that for any $x$, $\sum_{y\in{\mathsf{X}}}P(x,y)=1$; in other words, $P$ can be viewed as the transition probability of a Markov chain. Additionally, there is a probability distribution $\pi$ on ${\mathsf{X}}$ such that the weight matrix $P$ satisfies the detailed balance equation:
$\pi(x)P(x,y)=\pi(y)P(y,x)\,\,\forall x,y\in{\mathsf{X}};$
that is, $P$ is reversible. Hence, although graph signal processing techniques can be applied to any directed and weighted graph, in the remainder of this paper, we assume that the weight matrix $P$ corresponds to the transition probability of a reversible and irreducible Markov chain with respect to the unique stationary distribution $\pi$. Moreover, we endow $\mathbb{R}^{{\mathsf{X}}}$ with the inner product $\langle\cdot,\cdot\rangle_{\pi}$ introduced in the previous section. We direct the reader to the book [15] for the terminology and notation related to graph signal processing used in this paper.
In graph ${\mathcal{G}}$, for any vertex $x$, its out-degree is $d_{out}(x)=\sum_{x\rightarrow y}P(x,y)=1$. Consequently, the out-degree matrix for this graph is the identity matrix $\mathbb{Id}$. Therefore, the combinatorial graph Laplacian is defined as $L\coloneqq\mathbb{Id}-P$. Given that $\sigma(P)\subset[-1,1]$, it follows $\sigma(L)\subset[0,2]$.
Remark 1.
Given that $P$ satisfies the detailed balance equation with respect to $\pi$, we can define the following symmetric matrix:
$A(x,y)\coloneqq\pi(x)^{1/2}P(x,y)\pi(y)^{-1/2}\,\,\forall x,y\in{\mathsf{X}}.$
In matrix notation, if we let $D\coloneqq\mathop{\rm diag}(\pi)$ (i.e., the diagonal matrix with diagonal entries formed by the vector $\pi$), then, $A=D^{1/2}PD^{-1/2}$. Hence, $A$ and $P$ are similar matrices. Similar matrices represent the same linear operator in different bases, and they share all properties of the underlying operator, including the identical set of (real) eigenvalues. Consequently, we can use $A$ as a weight matrix instead of $P$. In this case, since $A$ is symmetric, we can treat our graph as a weighted undirected graph. However, a complication arises because the stationary distribution $\pi$ is either unknown or difficult to compute precisely, making it challenging to construct $A$ for use in graph signal processing.
Now, it is time to define signals on graphs in the node domain. Later, we will describe graph signals in the frequency domain.
Definition 2.
A graph signal $f$ is a mapping from the vertex set ${\mathsf{X}}$ to the reals.
In graph signal processing, the frequency of a signal is directly related to its variation across neighboring vertices. High signal variation indicates high frequency, while low variation suggests low frequency. To discuss these concepts accurately, we need to define the variation of a graph signal in our context. The next definition will provide this.
Definition 3.
A variation of a graph signal $f$ is defined as
$TV_{f}\coloneqq\frac{\left(\sum_{x\in{\mathsf{X}}}\sum_{x\rightarrow y}\pi(x)P(x,y)|f(x)-f(y)|^{2}\right)^{1/2}}{\left(\sum_{x\in{\mathsf{X}}}\pi(x)|f(x)|^{2}\right)^{1/2}}.$
In this definition, the denominator serves as a normalization constant to obtain a dimensionless quantity. The numerator quantifies the weighted difference of signals on vertices from their neighbors, with weights provided by the matrix $DP$. The value $\pi(x)P(x,y)$ represents how similar the vertices $x$ and $y$ are. Here, $\pi(x)$ indicates the importance of vertex $x$ in terms of the stationary distribution. If $\pi(x)\approx 0$, vertex $x$ holds little importance, as the number of visits to this vertex in the Markov chain is very low compared to other vertices according to the ergodic theorem. Additionally, $P(x,y)$ quantifies the (directed) similarity between vertex $x$ and $y$ or it can be interpreted as the importance of $y$ to $x$ because it is proportional to the number of transitions from vertex $x$ to $y$. Thus, the term $\pi(x)P(x,y)|f(x)-f(y)|^{2}$ captures the importance of vertex $x$ multiplied by the weighted variation of the graph signal $f$ from $x$ to $y$, with the significance of this variation determined by $P(x,y)$.
The following result is important for interpreting the eigenvalues of $L$ as frequencies.
Lemma 2.
We have $\sqrt{2\langle f,Lf\rangle_{\pi}}=TV_{f}$ for any graph signal $f$ with unit norm $\|f\|_{\pi}=1$.
Proof.
The proof of this lemma is very similar to [15, Remark 3.1].
For any $x\in{\mathsf{X}}$, we can write $Lf(x)=f(x)-\sum_{x\rightarrow y}P(x,y)f(y)=\sum_{x\rightarrow y}P(x,y)(f(x)-f(y))$ as $\sum_{x\rightarrow y}P(x,y)=1$. Hence,
$2\langle f,Lf\rangle_{\pi}$
$=\sum_{x\in{\mathsf{X}}}\sum_{x\rightarrow y}\pi(x)P(x,y)f(x)(f(x)-f(y))$
$\phantom{xxxxx}+\sum_{y\in{\mathsf{X}}}\sum_{y\rightarrow x}\pi(y)P(y,x)f(y)(f(y)-f(x))$
$=\sum_{x\in{\mathsf{X}}}\bigg{(}\sum_{x\rightarrow y}\pi(x)P(x,y)f(x)(f(x)-f(y))$
$\phantom{xxxxx}+\sum_{y\rightarrow x}\pi(y)P(y,x)f(y)(f(y)-f(x))\bigg{)}$
as $x\rightarrow y$ if and only if $y\rightarrow x$ by reversibility. Since $\pi(x)P(x,y)=\pi(y)P(y,x)$, we have
$\pi(y)P(y,x)f(y)=\pi(x)P(x,y)f(y).$
Hence,
$\sum_{x\rightarrow y}\pi(x)P(x,y)f(x)(f(x)-f(y))$
$\phantom{xxxxx}+\sum_{y\rightarrow x}\pi(y)P(y,x)f(y)(f(y)-f(x))$
$=\sum_{x\rightarrow y}\pi(x)P(x,y)(f(x)-f(y))^{2}.$
This implies that
$2\langle f,Lf\rangle_{\pi}=\sum_{x\in{\mathsf{X}}}\sum_{x\rightarrow y}\pi(x)P(x,y)(f(x)-f(y))^{2}=TV_{f}^{2}$
as $\|f\|_{\pi}=1$. This completes the proof. ∎
This result implies that $\langle f,Lf\rangle_{\pi}$ is high when the variation of $f$ is high, and low when the variation of $f$ is low. Therefore, we can use $\langle f,Lf\rangle_{\pi}$ to quantify the variation of graph signal $f$.
Let us write the eigenvalues of $L$ as $\{\lambda_{x}\}_{x\in{\mathsf{X}}}\subset[0,2]$, where one of them is zero, say $\lambda_{x_{*}}=0$. These eigenvalues have corresponding real eigenfunctions $\{f_{x}\}_{x\in{\mathsf{X}}}$, which form an orthonormal basis for the space $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle_{\pi})$ and $f_{x_{*}}={\bf 1}$. For any eigenfunction $f_{x}$, the following holds:
$\langle f_{x},Lf_{x}\rangle_{\pi}=\langle f_{x},\lambda_{x}f_{x}\rangle_{\pi}=\lambda_{x}.$
Therefore, if $\lambda_{x}$ is high, the variation of the corresponding eigenfunction $f_{x}$ is also high. This implies that eigenfunctions associated with high eigenvalues exhibit high frequencies. A similar conclusion can be drawn for eigenfunctions associated with low eigenvalues. The lowest frequency is $\lambda_{x_{*}}=0$, which corresponds to the constant eigenfunction ${\bf 1}$, where the variation is zero. Hence, we can view eigenvalues $\{\lambda_{x}\}_{x\in{\mathsf{X}}}$ of $L$ as frequencies.
Definition 4.
For any vertex $x\in{\mathsf{X}}$, its $k$-hop neighborhood, ${\mathcal{N}}_{k}(x)$, is the set all vertices that are part of a directed path with no more than $k$ edges starting from $x$. That is, $y\in{\mathcal{N}}_{k}(x)$ if there exists a directed path $\{x_{0},x_{1},\ldots,x_{l}\}$, where $l\leq k$, $x_{0}=x$, and $x_{l}=y$, such that $P(x_{i},x_{i+1})>0$ for all $i=0,\ldots,l-1$.
Now, it is time to give the definition of a graph filter.
Definition 5.
A (linear) graph filter $H$ is a linear operator on graph signals $f$ on ${\mathsf{X}}$.
In graph signal processing, the linear graph filter $H$ is considered to be local at $x$ if $Hf(x)$ only depends on $k$-hop neighbors of $x$, where $|{\mathcal{N}}_{k}(x)|\ll|{\mathsf{X}}|$. For instance,
$Pf(x)=\sum_{y\in{\mathsf{X}}}f(y)P(x,y)=\sum_{y\in{\mathcal{N}}_{1}(x)}f(y)P(x,y);$
that is, to compute $Pf(x)$, we need $1$-hop neighbors of $x$. Therefore, $P$ is a local graph filter. Similarly, $L$ is also a local graph filter. Indeed, $P$ and $L$ are called one-hop graph filters due to the following definition.
Definition 6.
A graph filter $H$ is called $k$-hop filter if $Hf(x)$ can be computed via the elements of ${\mathcal{N}}_{k}(x)$ for all $x\in{\mathsf{X}}$.
Let $Z$ represent a generic one-hop filter, such as $Z=P$, $Z=L$, or $Z=A$, among others. In graph signal processing, $Z$ serves as a generalization of the shift operator used in classical signal processing. For any positive integer $k$, $Z^{k}$ is a $k$-hop filter. Initially, our definition of a graph filter required only that the operator $H$ be linear, allowing any arbitrary linear operator to be chosen for $H$. However, we now focus on filters that can be expressed as polynomials of $Z$.
Definition 7.
Given a one-hop operator $Z$, the polynomial graph filter $H$ is a linear operator of the following form $H=p(Z)$ for some polynomial $p$ of some degree $K$. Hence,
$H=\sum_{k=0}^{K}a_{k}Z^{k},$
where $Z^{0}=\mathbb{Id}$.
To design a polynomial filter, one should first select a specific one-hop filter $Z$. In our case, this one-hop filter will be the combinatorial Laplacian $L$ since we have a precise interpretation of its eigenvalues as frequencies. Recall that the eigenvalues of $L$ are $\{\lambda_{x}\}_{x\in{\mathsf{X}}}\subset[0,2]$, where $\lambda_{x_{*}}=0$, with the corresponding real eigenfunctions $\{f_{x}\}_{x\in{\mathsf{X}}}$ forming an orthonormal basis for the space $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle{\pi})$ and $f_{x_{*}}=\mathbf{1}$. Thus, any graph signal $f$ has the following unique representation:
$f=\sum_{x\in{\mathsf{X}}}\langle f,f_{x}\rangle_{\pi}f_{x}.$
This representation leads to the definition of graph Fourier transform.
Definition 8.
[…]
Obviously $\mathcal{L}$ is a linear operator from $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle_{\pi})$ to $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle)$. Moreover, for any $x\in{\mathsf{X}}$, we have $\mathcal{L}(f_{x})=e_{x}$, where $e_{x}(y)\coloneqq 1_{\{x=y\}}$. Hence, $\mathcal{L}$ maps orthonormal basis $\{f_{x}\}_{x\in{\mathsf{X}}}$ of $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle_{\pi})$ to orhonormal basis $\{e_{x}\}_{x\in{\mathsf{X}}}$ of $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle)$. Therefore, it is a unitary transformation. Hence, the inverse of the GFT $\mathcal{L}$ is its adjoint: $\mathcal{L}^{-1}=\mathcal{L}^{*}$, where $\mathcal{L}^{*}$ is the unique operator from $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle)$ to $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle_{\pi})$ satisfying the following:
$\langle\mathcal{L}(f),{\hat{g}}\rangle=\langle f,\mathcal{L}^{*}({\hat{g}})\rangle_{\pi}\,\,\forall f\in(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle_{\pi}),{\hat{g}}\in(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle).$
Hence $\mathcal{L}^{-1}({\hat{g}})(x)=\sum_{y\in{\mathsf{X}}}f_{y}(x){\hat{g}}(y)$.
Let $H$ be a polynomial graph filter of degree $K$; that is, $H=\sum_{k=0}^{K}a_{k}L^{k}\eqqcolon p(L)$. This filter then acts on any graph signal $f$ with GFT ${\hat{f}}$ in the following manner:
$H(f)=\sum_{x\in{\mathsf{X}}}{\hat{f}}(x)\,p(\lambda_{x})\,f_{x}.$
Hence, in the frequency domain, the polynomial filter $H$ can be described as follows:
${\hat{H}}\coloneqq\mathop{\rm diag}(\{p(\lambda_{x})\}_{x\in{\mathsf{X}}})=p(\Lambda),$
where $\Lambda\coloneqq\mathop{\rm diag}(\{\lambda_{x}\}_{x\in{\mathsf{X}}})$; that is, $\Lambda$ is the diagonal matrix containing the eigenvalues of $L$ as its diagonal entries. Given ${\hat{H}}$, we can describe the action of the polynomial filter $H$ in the frequency domain as follows:
${\hat{H}}({\hat{f}})=\mathop{\rm diag}(\{{\hat{f}}(x)\,p(\lambda_{x})\}_{x\in{\mathsf{X}}})=\mathcal{L}(H(f)).$
Unfortunately, since $\pi$ is generally unavailable, it is in general unfeasible to compute the GFT of any graph signal $f$ and frequency representation of any polynomial graph filter $H$. Consequently, filtering operations must be performed in the node domain instead of the frequency domain. Nevertheless, we can still gain insight into designing graph filters in the node domain by examining their effects in the frequency domain.
Now, we provide an informal definition of the problem we aim to solve in this paper. A precise mathematical description will follow later.
Informal Definition of the Problem: For each degree $K$, find the optimal polynomial filter $H_{K}$ with the corresponding polynomial $p_{K}$ such that $H_{K}(f)$ converges to $\pi(f){\bf 1}$ as quickly as possible. To achieve this convergence, the polynomial $p_{K}$ should attenuate the eigenvalues (or frequencies) other than 00 of $L$. Therefore, the polynomial graph filter should function as a low-pass filter.
Revisiting Ergodic Theorem
Note that the iteration in the ergodic theorem can be written as a polynomial of $P$ in the following form:
$\frac{(\mathbb{Id}+P+\ldots+P^{t-1})}{t}\eqqcolon p_{t}(P),$
where
$p_{t}(z)\coloneqq\sum_{k=0}^{t-1}\frac{1}{t}z^{t}=\frac{1}{t}\frac{1-z^{t}}{1-z},\,\,\,z\in[-1,1).$
If we define the following polynomial $q_{t}(z)\coloneqq p_{t}(1-z)$, then we can write the same expression in terms of $L$ as follows:
$p_{t}(P)$ $=q_{t}(L)=\frac{1}{t}\frac{1-(1-L)^{t}}{1-(1-L)}$
$=\frac{1}{t}\frac{1-\sum_{k=0}^{t}{t\choose k}(-L)^{k}}{L}$
$=\sum_{k=1}^{t}\frac{1}{t}{t\choose k}(-1)^{k-1}L^{k-1}$
$\eqqcolon\sum_{k=0}^{t-1}a_{e}(k)\,L^{k}.$
If we define the polynomial graph filter $H_{t}\coloneqq q_{t}(L)$, then we observe that we are indeed applying the polynomial graph filter $H_{t}$ to the graph signal $f$ in each iteration of the ergodic theorem. As $t$ approaches infinity, this graph filter attenuates the frequencies different from zero. Specifically,
$q_{t}(z)\rightarrow 0\,\,\text{as}\,\,t\rightarrow\infty$
for $z\in(0,2]$. This can be proved very easily since $p_{t}(z)\rightarrow 0$ as $t\rightarrow\infty$ for $z\in[-1,1)$. Consequently, $H_{t}(f)\rightarrow\pi(f){\bf 1}$. Hence, in each iteration of the ergodic theorem, we are applying a low-pass filter to the graph signal $f$. However, this low-pass filter is not the optimal one. The goal of this paper is to find or approximate the optimal low-pass filter in each iteration. $\square$
Given the non-zero eigenvalues $\{\lambda\}_{x\in{\mathsf{X}}\setminus\{x_{*}\}}\subset(0,2]$ of $L$, it is not always the case that $q_{t}(\lambda_{x})=0$ for all $x\in{\mathsf{X}}\setminus\{x_{*}\}$ for sufficiently large $t$ values, where $q_{t}$ is the polynomial used in the ergodic theorem. Indeed, if we have complete knowledge of the eigenvalues $\{\lambda\}_{x\in{\mathsf{X}}}$ of $L$, then we can use a Lagrange polynomial $q$ of degree $|{\mathsf{X}}|-1$ [16, Definition 6.1] to achieve:
$q(\lambda_{x_{*}})=1\,\,\text{and}\,\,q(\lambda_{x})=0\,\,\text{for}\,\,x\neq x_{*}.$
In this case, we set $q_{t}^{L}(z)=(1+z+\ldots+z^{t-|{\mathsf{X}}|})\,q(z)$ for $t\geq|{\mathsf{X}}|$. Hence
$q_{t}^{L}(L)\,f=\sum_{x\in{\mathsf{X}}}{\hat{f}}(x)\,q_{t}^{L}(\lambda_{x})\,f_{x}=\pi(f){\bf 1}.$
However, in general, we do not have access to the full set of eigenvalues of $L$ and so constructing $q_{t}^{L}$ is unfeasible. In general, the most that one can do is to estimate some important eigenvalues different than $\lambda_{x_{*}}$. For instance, we may obtain some lower bound to the second smallest eigenvalue of $L$ [17] [18, Section 9.2]. Indeed, if $\lambda_{x_{**}}$ denotes the second smallest eigenvalue of $L$, then $\lambda_{x_{**}}$ is equal to the spectral gap of $P$. In this case, if we have some lower bound on this second smallest eigenvalue:
$\lambda_{x_{**}}\geq\lambda_{{\mathsf{low}}},$
then the best approach would be to find a polynomial that satisfies:
$q_{t}(\lambda_{x_{*}})=1\,\,\text{and}\,\,q_{t}(\lambda)\approx 0\,\,\text{for}\,\,2\geq\lambda\geq\lambda_{{\mathsf{low}}}.$
Here, we want to converge to zero for values in $[\lambda_{{\mathsf{low}}},2]$ as quickly as possible through polynomials of the form $q_{t}(z)=\sum_{k=0}^{t-1}a(k)z^{k}$ while maintaining $q_{t}(\lambda_{x_{*}})=1$. Thus, we can interpret our problem as designing an optimal (in some sense) polynomial low-pass graph filter with one-hop operator $L$ to filter out frequencies greater than $\lambda_{{\mathsf{low}}}$. In the following, we assume that $\lambda_{{\mathsf{low}}}$ is fixed. Using this assumption, we provide a precise mathematical description of the optimization problem we aim to solve.
Let $P_{K}[\lambda_{{\mathsf{low}}},2]$ be the set of all polynomials of degree at most $K$. Then, at each time step $t\geq 2$, we want to solve the following optimization problem:
${\bf(OPT_{t})}\min_{p\in P_{t-1}[\lambda_{{\mathsf{low}}},2]}\|p\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}\,\,\text{subject to}\,\,p(0)=1.$
Here
$\|p\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}\coloneqq\sup_{z\in[\lambda_{{\mathsf{low}}},2]}|p(z)|.$
Let us now give a motivation for this optimization problem. Note that, originally, we want to minimize $\|p_{t}(L)f-\pi(f){\bf 1}\|_{\pi}$ for any unit norm graph signal $f$, where $p_{t}\in P_{t-1}[\lambda_{{\mathsf{low}}},2]$ with $p_{t}(0)=1$, for each time step $t\geq 2$ as any graph signal is viewed as an element of the inner product space $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle_{\pi})$ [^1] In this context, we can use any norm we prefer because the interpretation at the end will be the same.. An upper bound to this term can be obtained as follows:
$\|p_{t}(L)f-\pi(f){\bf 1}\|_{\pi}$ $=\left\|\sum_{x\in{\mathsf{X}}}{\hat{f}}(x)p_{t}(\lambda_{x})f_{x}-{\hat{f}}(x_{*})f_{x_{*}}\right\|_{\pi}$
$=\left\|\sum_{x\neq x_{*}}{\hat{f}}(x)p_{t}(\lambda_{x})f_{x}\right\|_{\pi}$
$\leq\sum_{x\neq x_{*}}|{\hat{f}}(x)||p_{t}(\lambda_{x})|\left\|f_{x}\right\|_{\pi}$
$=\sum_{x\neq x_{*}}|{\hat{f}}(x)||p_{t}(\lambda_{x})|$
$\leq\left(\sum_{x\neq x_{*}}|{\hat{f}}(x)|\right)\sup_{x\neq x_{*}}|p_{t}(\lambda_{x})|$
In the last inequality above, the first term is fixed. Thus, our objective should be to minimize the second term by choosing $p_{t}$ appropriately. However, because we do not have access to the full eigenvalue structure, we cannot directly minimize the second term. Given the lower bound $\lambda_{{\mathsf{low}}}$ for the second smallest eigenvalue of $L$, we can instead try to minimize $\|p_{t}\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}$, which leads to the optimization problem ${\bf(OPT_{t})}$.
IV Filter Design with Bernstein Polynomials
For each time step $t\geq 2$, we want to design a polynomial $p$ of degree at most $t-1$ so that
$p(\lambda_{x_{*}})=1\,\,\text{and}\,\,p(z)\approx 0\,\,\text{for}\,\,2\geq z\geq\lambda_{{\mathsf{low}}}.$
In this sub-section, instead of pursuing the optimal design formulated in ${\bf(OPT_{t})}$, we approach the problem differently. Firstly, the behavior of the polynomial $p$ over the interval $(0,\lambda_{{\mathsf{low}}})$ is not critical. Therefore, let us fix a continuous function $g$ on $[0,2]$ such that $g(\lambda_{x_{*}})=g(0)=1$ and $g(z)=0$ for $2\geq z\geq\lambda_{{\mathsf{low}}}$. We have the freedom to choose the behavior of $g$ on the interval $(0,\lambda_{{\mathsf{low}}})$ as desired. This choice is also part of the design process. Note that Weierstrass approximation theorem [19] states that any continuous function over compact domain can be approximated via polynomials under sup-norm. This theorem is generally proved constructively using Bernstein polynomials (see the proof of [19, Theorem 1.1]). Therefore, we can use Bernstein polynomials to approximate our function $g$ over $[0,2]$. To this end, let us first define Bernstein polynomials.
Definition 9.
Given any function $h$ that is bounded on $[0,1]$, we define its Bernstein polynomial of degree $K$ by
$B_{K}(h;z)\coloneqq\sum_{l=0}^{K}h\left(\frac{l}{K}\right){K\choose l}z^{l}\,(1-z)^{K-l}.$
One can prove that (see [19, Theorem 1.2])
$\|h-B_{K}(h)\|_{\infty}^{[0,1]}\leq\frac{3}{2}\,\omega_{h}\left(\frac{1}{\sqrt{K}}\right),$
where $\omega_{h}$ is the modulus of continuity of $h$. Since our function $g$ is defined over $[0,2]$ and Bernstein polynomials are defined over $[0,1]$, we need to translate this discussion into the interval $[0,2]$. The Bernstein polynomial of degree $K$ of $g$ is defined as
$B_{K}(g;z)\coloneqq\sum_{l=0}^{K}g\left(\frac{2l}{K}\right){K\choose l}\left(\frac{z}{2}\right)^{l}\,\left(1-\frac{z}{2}\right)^{K-l}.$
Then, by above result, we have
$\|g-B_{K}(g)\|_{\infty}^{[0,2]}\leq\frac{3}{2}\,\omega_{g}\left(\frac{2}{\sqrt{K}}\right).$
Additionally, $B_{K}(g;\lambda_{x_{*}})=B_{K}(g;0)=1$. Thus, $B_{K}(g)$ is a strong candidate for our polynomial filter. However, given the above rate of convergence result, to achieve a sharp decrease in the function value of $B_{K}(g)$ over the interval $[\lambda_{{\mathsf{low}}},2]$, it makes sense to select $g$ or adjust its behavior on the interval $(0,\lambda_{{\mathsf{low}}})$ so that $\omega_{g}$ is minimized at $\frac{2}{\sqrt{K}}$. Consequently, we propose the following optimization problem for each time step $t\geq 2$:
${\bf(OPT^{B}_{t})}$ $\min_{g\in C[0,2]}\omega_{g}\left(\frac{2}{\sqrt{t-1}}\right)\,\,$ $\text{subject to}\,\,g(0)=1,\,\,g([\lambda_{{\mathsf{low}}},2])=0.$
Solving this optimization problem for any $t\geq 2$ is generally challenging and impractical to perform at each time step $t\geq 2$. However, as $t$ approaches infinity, we can demonstrate that the triangle function $g_{\Delta}$, which decreases from 1 to zero as we move from 0 to $\lambda_{{\mathsf{low}}}$, is proved to be asymptotically optimal for this problem. Therefore, for each $t\geq 2$, we can select $g$ as $g_{\Delta}$ to design our Bernstein filter.
Asymptotically Optimal Bernstein Filter $B_{K}(g_{\Delta};z)\coloneqq\sum_{l=0}^{K}g_{\Delta}\left(\frac{2l}{K}\right){K\choose l}\left(\frac{z}{2}\right)^{l}\,\left(1-\frac{z}{2}\right)^{K-l}$
Let us now demonstrate that the triangle function is asymptotically optimal. First, the modulus of continuity, denoted as $\omega_{\Delta}$, for the triangle function $g_{\Delta}$ is given by the following:
$\omega_{\Delta}(\delta)=\begin{cases}m\delta&\text{if}\,\,\delta\in[0,\lambda_{{\mathsf{low}}})\\ 1&\text{otherwise}\end{cases}$
where $m\coloneqq 1/\lambda_{{\mathsf{low}}}$. Note that if $2/\sqrt{t-1}\geq\lambda_{{\mathsf{low}}}$, then for any $g\in C[0,2]$ satisfying the constraints in ${\bf(OPT^{B}_{t})}$, we have $\omega_{g}\left(2/\sqrt{t-1}\right)=1$. Hence, any feasible $g\in C[0,2]$ solves ${\bf(OPT^{B}_{t})}$. For this reason, in the following, we assume that $2/\sqrt{t-1}<\lambda_{{\mathsf{low}}}$.
Proposition 1.
If $g\in C[0,2]$ is an optimal solution of ${\bf(OPT^{B}_{t})}$, where $r\coloneqq 2/\sqrt{t-1}<\lambda_{{\mathsf{low}}}$, then
$\omega_{\Delta}(r)\frac{n+\lambda}{n+1+\lambda}\leq\omega_{g}(r)<\omega_{\Delta}(r).$
Here, $\lambda_{{\mathsf{low}}}=nr+\lambda r$ and $\lambda\in(0,1)$. Hence, as $t\rightarrow\infty$, we have $\lambda\rightarrow 0$, $n\rightarrow\infty$, and $\frac{n+\lambda}{n+1+\lambda}\rightarrow 1$. This implies that $g_{\Delta}$ is asymptotically optimal.
Proof.
Let $g\in C[0,2]$ be an optimal solution [^2] If the optimal solution does not exist, we can use an $\varepsilon$-optimal solution for a sufficiently small $\varepsilon>0$. In this case, we should replace $\omega_{\Delta}$ with $\omega_{\Delta}+\varepsilon$ in the second inequality stated in the proposition. of ${\bf(OPT^{B}_{t})}$, where $r<\lambda_{{\mathsf{low}}}$. By [19, Lemma 1.3], we have
11 11 $=\omega_{g}(nr+\lambda r)\leq(n+\lambda+1)\omega_{g}(r)$
$\leq(n+\lambda+1)\omega_{\Delta}(r).$
But note that $1=\omega_{\Delta}(nr+\lambda r)=(n+\lambda)\omega_{\Delta}(r)$. Hence
$(n+\lambda)\omega_{\Delta}(r)\leq(n+\lambda+1)\omega_{g}(r)\leq(n+\lambda+1)\omega_{\Delta}(r).$
Dividing both sides with $(n+\lambda+1)$ establishes the result. ∎
Another justification for why the triangle function is a good candidate is the following. Instead of considering the space $C[0,2]$, we can work with the set of continuously differentiable, except possibly at $\lambda_{{\mathsf{low}}}$, functions on $[0,2]$, denoted as $C^{1}[0,2]$. Since estimating the modulus of continuity exactly can be challenging, we can use the inequality $\omega_{g}(\delta)\leq\|g^{(1)}\|_{\infty}^{[0,2]}\delta$ for all $\delta\geq 0$ to obtain an upper bound on the modulus of continuity. This allows us to state a more manageable optimization problem, which does not depend on time.
${\bf(\widehat{OPT}_{B})}$ $\min_{g\in C^{1}[0,2]}\|g^{(1)}\|_{\infty}^{[0,2]}\,\,$ $\text{subject to}\,\,g(0)=1,\,\,g([\lambda_{{\mathsf{low}}},2])=0.$
The solution of the last problem is again the triangle function $g_{\Delta}$ as stated in the next result.
Proposition 2.
The triangle function $g_{\Delta}$ is an optimal solution of ${\bf(\widehat{OPT}_{B})}$.
Proof.
According to the mean value theorem, for any function $g\in C^{1}[0,2]$ satisfying $g(0)=1$ and $g([\lambda_{{\mathsf{low}}},2])=0$, there exists an $l\in(0,\lambda_{{\mathsf{low}}})$ such that $|g^{(1)}(l)|=1/\lambda_{{\mathsf{low}}}$, and so, $\|g^{(1)}\|_{\infty}^{[0,2]}\geq 1/\lambda_{{\mathsf{low}}}$. Furthermore, note that for the triangle function, $\|g^{(1)}_{\Delta}\|_{\infty}^{[0,2]}=1/\lambda_{{\mathsf{low}}}$. $\square$
V Filter Design with Chebyshev Polynomials
Recall that for each time step $t\geq 2$, we want to solve the following optimization problem:
${\bf(OPT_{t})}\min_{p\in P_{t-1}[\lambda_{{\mathsf{low}}},2]}\|p\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}\,\,\text{subject to}\,\,p(0)=1.$
In this section, we establish that the optimal solution of ${\bf(OPT_{t})}$ is a properly normalized Chebyshev polynomial of degree $t-1$, for any $t\geq 2$. Since we have explicit forms of Chebyshev polynomials of any degree, we can easily implement the corresponding filter. Before proving this result, let us briefly discuss Chebyshev polynomials and outline one of their key properties, which will lead us to the optimal solution of ${\bf(OPT_{t})}$. This property can be stated as follows: among all polynomials with a fixed degree and a sup-norm less than $1$ over any interval $[a,b]$, Chebyshev polynomials are the ones that increase as quickly as possible compared to other polynomials outside this interval. Using this property, it becomes clear that Chebyshev polynomials are the optimal solutions of ${\bf(OPT_{t})}$.
Chebyshev polynomials are originally defined over the interval $[-1,1]$. Therefore, we begin by defining these polynomials over this interval and stating their maximum growth property. Then, we translate everything to the interval $[\lambda_{{\mathsf{low}}},2]$ in our problem formulation.
Definition 10.
The Chebyshev polynomial of degree $K$ is defined as follows:
$T_{K}(z)\coloneqq\cos(K\theta),$
where $z=\cos(\theta)$ and $\theta\in[0,\pi]$.
Note that since $\cos(K\theta)$ is a polynomial of $\cos(\theta)$ with degree $K$ [19, Exercise 1.6, p. 43], $T_{K}$ is well-defined. Moreover, since it is a polynomial, once we know its coefficients, we know it as a function for all $z$. The following result states that Chebyshev polynomials grows as rapidly as possible outside of $[-1,1]$.
Theorem 2.
[19, Theorem 1.10, p. 31] If $p\in P_{K}[-1,1]$ and $|z_{0}|\geq 1$, then
$|p^{(l)}(z_{0})|\leq\|p\|_{\infty}^{[-1,1]}\,|T_{K}^{(l)}(z_{0})|,\,\,l=0,1,\ldots,K,$
and in particular, we have
$|p(z_{0})|\leq\|p\|_{\infty}^{[-1,1]}\,|T_{K}(z_{0})|.$
Using above maximal growth condition, we can now prove the main result of this section. Before doing that, let us define the Chebyshev polynomial of degree $K$ over the interval $[\lambda_{{\mathsf{low}}},2]$ using a linear transformation as follows:
$T_{K}^{[\lambda_{{\mathsf{low}}},2]}(z)\coloneqq T_{K}\left(\frac{2z-2-\lambda_{{\mathsf{low}}}}{2-\lambda_{{\mathsf{low}}}}\right).$
Hence, $T_{K}^{[\lambda_{{\mathsf{low}}},2]}(\lambda_{{\mathsf{low}}})=T_{K}(-1)$ and $T_{K}^{[\lambda_{{\mathsf{low}}},2]}(2)=T_{K}(1)$. Note that $T_{K}^{[\lambda_{{\mathsf{low}}},2]}(0)\neq 1$, and so, $T_{K}^{[\lambda_{{\mathsf{low}}},2]}$ is not a feasible point for ${\bf(OPT_{t})}$ in this form. Hence, we need to normalize it as follows:
$p_{K}^{*}(z)\coloneqq\frac{T_{K}^{[\lambda_{{\mathsf{low}}},2]}(z)}{T_{K}^{[\lambda_{{\mathsf{low}}},2]}(0)}.$
Now, we have $p_{K}^{*}(0)=1$. Moreover, we also have
$\|p_{K}^{*}\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}=\frac{\left\|T_{K}^{[\lambda_{{\mathsf{low}}},2]}\right\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}}{\left|T_{K}^{[\lambda_{{\mathsf{low}}},2]}(0)\right|}=\frac{1}{\left|T_{K}^{[\lambda_{{\mathsf{low}}},2]}(0)\right|}$
Theorem 3.
We have $p_{t-1}^{*}\in\mathop{\rm arg\,min}{\bf(OPT_{t})}$ for any $t\geq 2$. Hence, the optimal polynomial filter is given by properly normalized Chebyshev polynomial of degree $t-1$ at each time step $t\geq 2$.
Proof.
Fix any $t\geq 2$. Suppose that there exists $p\in P_{t-1}[\lambda_{{\mathsf{low}}},2]$ such that $p(0)=1$ and $\|p\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}<\|p_{t-1}^{*}\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}$. Our goal is to reach a contradiction. Note that Theorem 2 is still true for Chebyshev polynomials over the interval $[\lambda_{{\mathsf{low}}},2]$. Therefore, as $0\notin[\lambda_{{\mathsf{low}}},2]$, we have
$|p(0)|\leq\|p\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}\left|T_{K}^{[\lambda_{{\mathsf{low}}},2]}(0)\right|.$
But this implies that
$|p(0)|<\|p_{t-1}^{*}\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}\left|T_{K}^{[\lambda_{{\mathsf{low}}},2]}(0)\right|=1$
as
$\|p_{t-1}^{*}\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}=\frac{1}{\left|T_{K}^{[\lambda_{{\mathsf{low}}},2]}(0)\right|}.$
This is a contradiction as we assumed $p(0)=1$. ∎
Note that Chebyshev polynomials over $[-1,1]$ satisfy the following recursive relation [19, p. 55]:
$T_{K+1}(z)=2z\,T_{K}(z)-T_{K-1}(z)\,\,K\geq 1.$
Hence we have
$T_{K+1}^{[\lambda_{{\mathsf{low}}},2]}(z)=\frac{2z-2-\lambda_{{\mathsf{low}}}}{2-\lambda_{{\mathsf{low}}}}\,T_{K}^{[\lambda_{{\mathsf{low}}},2]}(z)-T_{K-1}^{[\lambda_{{\mathsf{low}}},2]}(z).$
For each $K\geq 1$, let us define
$\alpha_{K}\coloneqq\frac{T_{K}^{[\lambda_{{\mathsf{low}}},2]}(0)}{T_{K+1}^{[\lambda_{{\mathsf{low}}},2]}(0)},\,\,\beta_{K}\coloneqq\frac{T_{K-1}^{[\lambda_{{\mathsf{low}}},2]}(0)}{T_{K+1}^{[\lambda_{{\mathsf{low}}},2]}(0)}.$
Then, we have the following recursive relation between optimal Chebyshev filters:
Recursive Implementation of Chebyshev Filter $p_{K+1}^{*}(L)=\alpha_{K}\,\frac{2L-2-\lambda_{{\mathsf{low}}}}{2-\lambda_{{\mathsf{low}}}}p_{K}^{*}(L)-\beta_{K}\,p_{K-1}^{*}(L)$
This formula enables us to calculate the output of the optimal filter recursively, significantly reducing the amount of algebraic computation required in each iteration, as it eliminates the need to compute $p_{K+1}^{*}(L)f$ from scratch. By storing $p_{K}^{*}(L)f$ and $p_{K-1}^{*}(L)f$ in memory, we can easily calculate $p_{K+1}^{*}(L)f$ using the recursive formula.
VI Filter Design with Legendre Polynomials
In this section, we reformulate the optimization problem ${\bf(OPT_{t})}$ by using the $l_{2}$-norm instead of the sup-norm. It turns out that the optimal solution is weighted combination of properly scaled and normalized Legendre polynomials.
For each time step $t\geq 2$, our goal is to solve the following optimization problem:
${\bf(\widehat{OPT}_{t})}\min_{p\in P_{t-1}[\lambda_{{\mathsf{low}}},2]}\|p\|_{2}^{[\lambda_{{\mathsf{low}}},2]}\,\,\text{subject to}\,\,p(0)=1.$
Here, the $l_{2}$-norm is with respect to the Lebesgue measure over $[\lambda_{{\mathsf{low}}},2]$:
$\|p\|_{2}^{[\lambda_{{\mathsf{low}}},2]}\coloneqq\left(\int_{[\lambda_{{\mathsf{low}}},2]}|p(x)|^{2}\,m(dx)\right)^{1/2}.$
In general, for any Lebesgue integrable weight function $w:[\lambda_{{\mathsf{low}}},2]\rightarrow(0,\infty)$, we can define the weighted $l_{2}$-norm as
$\|p\|_{2,w}^{[\lambda_{{\mathsf{low}}},2]}\coloneqq\left(\int_{[\lambda_{{\mathsf{low}}},2]}|p(x)|^{2}\,w(x)\,m(dx)\right)^{1/2}.$
Therefore, depending on $w$, the solution of the optimization problem $\mathbf{(\widehat{OPT}_{t})}$ varies. However, because we lack knowledge of the eigenvalues of $P$ in the interval $[\lambda_{{\mathsf{low}}},2]$, we do not know how to assign weights to the points within this interval. Thus, it is more reasonable to assign equal weights, i.e., $w(x)=1$, to all the points in $[\lambda_{{\mathsf{low}}},2]$, which leads to the classical $l_{2}$-norm.
To solve the above optimization problem, we need to define orthonormal polynomials over $[\lambda_{{\mathsf{low}}},2]$ with respect to the classical $l_{2}$-norm. Note that if our interval is $[-1,1]$, then Legendre polynomials $\{L_{n}\}$ are the orthogonal polynomials [19, p. 53], where these polynomials are recursively defined as follows:
$L_{0}(z)=1,\,\,L_{1}(z)=z,\,\,$
$(n+1)L_{n+1}(z)=(2n+1)zL_{n}(z)-nL_{n-1}(z)\,\,n\geq 1.$
Note that for each $n\geq 0$, we have [19, Exercise 2.11, p. 62]
$\int_{[-1,1]}L_{n}(z)^{2}m(dz)=\frac{2}{2n+1}.$
Hence, to obtain an orthonormal system, we need to normalize the Legendre polynomials as follows:
$\hat{L}_{n}(z)\coloneqq\sqrt{n+\frac{1}{2}}\,L_{n}.$
These normalized Legendre polynomials satisfy the following recursive formula:
$\sqrt{\frac{2(n+1)^{2}}{2n+3}}\hat{L}_{n+1}(z)=\sqrt{2(2n+1)}z\hat{L}_{n}(z)-\sqrt{\frac{2n^{2}}{2n-1}}\hat{L}_{n-1}(z).$
Now, we can transform the orthonormality properties of these polynomials to the interval $[\lambda_{{\mathsf{low}}},2]$ by simply applying a linear transformation, which leads to the following orthonormal (with respect to the classical $l_{2}$-norm) polynomials over $[\lambda_{{\mathsf{low}}},2]$:
$\tilde{L}_{n}(z)\coloneqq\frac{2}{2-\lambda_{{\mathsf{low}}}}\,\hat{L}_{n}\left(\frac{2z-2-\lambda_{{\mathsf{low}}}}{2-\lambda_{{\mathsf{low}}}}\right)\,\,n\geq 0.$
Therefore, they satisfy the following recursive formula:
$\sqrt{\frac{2(n+1)^{2}}{2n+3}}\tilde{L}_{n+1}(z)$
$=\sqrt{2(2n+1)}\,\frac{2z-2-\lambda_{{\mathsf{low}}}}{2-\lambda_{{\mathsf{low}}}}\,\tilde{L}_{n}(z)-\sqrt{\frac{2n^{2}}{2n-1}}\tilde{L}_{n-1}(z).$
If we define
$\alpha_{n}\coloneqq\sqrt{\frac{2n^{2}}{2n-1}},\,\,\beta_{n}\coloneqq\sqrt{2(2n+1)}\,\,n\geq 1,$
then we can write the above recursive formula as
$\alpha_{n+1}\tilde{L}_{n+1}(z)=\beta_{n}\,\frac{2z-2-\lambda_{{\mathsf{low}}}}{2-\lambda_{{\mathsf{low}}}}\,\tilde{L}_{n}(z)-\alpha_{n}\tilde{L}_{n-1}(z).$
Now, we present the main result of this section. To do so, for each $t\geq 2$ and $k\geq 0$, we define
$\xi_{k}^{t-1}\coloneqq\frac{\tilde{L}_{k}(0)}{\sum_{k=0}^{t-1}\tilde{L}_{k}(0)^{2}}.$
Using these coefficients, for each $t\geq 2$, let us also define the following polynomial:
$p_{t-1}^{*}(z)\coloneqq\sum_{k=0}^{t-1}\xi_{k}^{t-1}\,\tilde{L}_{k}(z).$
Theorem 4.
We have $p_{t-1}^{*}\in\mathop{\rm arg\,min}{\bf(\widehat{OPT}_{t})}$ for any $t\geq 2$. Hence, the optimal polynomial filter is given by weighted combination of properly normalized and scaled Legendre polynomials of degree at most $t-1$, at each time step $t\geq 2$.
Proof.
The proof of this problem is quite classical and can be found in any book on orthogonal polynomials. We refer readers to [20, p. 78]. For completeness, we provide the full proof here, which is quite elementary.
Note that any polynomial $p\in P_{t-1}[\lambda_{{\mathsf{low}}},2]$ can be written
$p=\sum_{k=0}^{t-1}c_{k}\,\tilde{L}_{k}.$
If $p$ is a feasible point of ${\bf(\widehat{OPT}_{t})}$, then
$p(0)=\sum_{k=0}^{t-1}c_{k}\,\tilde{L}_{k}(0)=1.$
Hence, by Cauchy’s inequality, we have
$1\leq\left(\sum_{k=0}^{t-1}c_{k}^{2}\right)\left(\sum_{k=0}^{t-1}\tilde{L}_{k}(0)^{2}\right).$
Therefore, for any feasible $p$, we have
$\left(\|p\|_{2}^{[\lambda_{{\mathsf{low}}},2]}\right)^{2}=\sum_{k=0}^{t-1}c_{k}^{2}\geq\frac{1}{\left(\sum_{k=0}^{t-1}\tilde{L}_{k}(0)^{2}\right)}.$
Note that we have equality above if and only if
$c_{k}=\frac{\tilde{L}_{k}(0)}{\sum_{k=0}^{t-1}\tilde{L}_{k}(0)^{2}}\eqqcolon\xi_{k}^{t-1}.$
This completes the proof. ∎
Note that we can implement this optimal filter recursively at two different time scales. Specifically, for each $K\geq 0$, we define
$\gamma_{K}\coloneqq\frac{\sum_{k=0}^{K}\tilde{L}_{k}(0)^{2}}{\sum_{k=0}^{K+1}\tilde{L}_{k}(0)^{2}}.$
Then the optimal filter can be computed via the following coupled iterations:
Recursive Implementation of Legendre Filter $p_{K+1}^{*}(L)=\gamma_{K}\,p_{K}^{*}(L)+\xi_{K+1}^{K+1}\tilde{L}_{K+1}(L)$ $\tilde{L}_{K+1}(L)=\frac{\beta_{K}}{\alpha_{K+1}}\frac{2L-2-\lambda_{{\mathsf{low}}}}{2-\lambda_{{\mathsf{low}}}}\tilde{L}_{K}(L)-\frac{\alpha_{K}}{\alpha_{K+1}}\tilde{L}_{K-1}(L)$
This formula allows us to compute the output of the optimal filter recursively, greatly reducing the algebraic computation needed in each iteration, as it avoids recomputing $p_{K+1}^{*}(L)f$ from the beginning. By storing $p_{K}^{*}(L)f$, $\tilde{L}_{K}(L)f$, and $\tilde{L}_{K-1}(L)f$ in memory, we can efficiently compute $p_{K+1}^{*}(L)f$ using the coupled recursive formulas.
VII Numerical Experiments
In this section, we numerically test the performance of the optimal filters derived in the previous sections. Our goal is to compare these filters with the iterative process described in the classical ergodic theorem. To achieve this, we consider two specific examples, both of which are detailed in Section II: the random walk on graphs and the Glauber chains.
First, we examine the random walk on graphs. Next, we turn our attention to the Glauber chains. By comparing the outcomes from these two examples, we gain a deeper understanding of the strengths and limitations of the optimal filters. This comparison also provides insights into the practical applications of these filters in various Markov chains.
VII-A Random Walk on a Graph
In this section, we consider a very simple random walk on a cycle [17, Example 2.1]. Let $p$ be an odd number. We place the integers $\mathbb{Z}_{p}\coloneqq\{0,1,2,\ldots,p-1\}$ on a circle in the given order and connect each consecutive point in the circle. This forms our graph. In Figure 1, an example is given where $p=11$.
012345678910
Figure 1: A circle with $p$ integers placed evenly around it, where $p$ is an odd number.
We consider a random walk on this graph. Therefore, we have
$P(x,y)=\begin{cases}\frac{1}{2}&\text{if}\,\,|x-y|\in\{1,p-1\}\mod p\\ 0&\,\,\text{otherwise}\end{cases}$
and $\pi(x)=1/p$. For this Markov chain, by [17, Corollary 1] (see also [17, p. 44]), one can obtain the following lower bound to the second smallest eigenvalue $\lambda_{x_{**}}$ of $L$ as follows:
$\lambda_{x_{**}}\geq\lambda_{{\mathsf{low}}}\coloneqq\frac{8p}{(p-1)^{2}(p+1)}.$
Indeed, for this random walk, it is possible to compute the exact eigenvalues of $L$ for small $p$ values. However, since this is a toy example in which we want to demonstrate the effectiveness of our method, and computing the eigenvalues of $L$ for other, more complicated models is not feasible, we use the above estimate instead of the exact value.
In the numerical experiment, we take $p=11$ and generate a random function $f$ as follows:
$f=\begin{bmatrix}8.53,6.22,3.50,5.13,4.01,0.75,2.39,1.23,1.83,2.39,4.17\end{bmatrix}$
In this case, $\lambda_{{\mathsf{low}}}=0.0733$. We apply the ergodic average, Bernstein polynomial filter, Chebyshev polynomial filter, and Legendre polynomial filter up to degree $20$, and obtain the results depicted in Figure 2.
Figure 2: A maximum absolute error as a function of degree of the filter.
In this figure, the maximum absolute error is computed as
$\max_{x\in{\mathsf{X}}}\bigg{|}p(L)f(x)-\sum_{x\in{\mathsf{X}}}f(x)\,\pi(x)\bigg{|}$
for any polynomial filter $p(L)$. This gives the worst deviation of the filter output from the desired value with respect to the initial position. As one can see, the Bernstein polynomial filter is slightly better than the ergodic average. However, the Chebyshev and Legendre polynomial filters outperform the ergodic average significantly, as expected. Indeed, for these polynomials, the maximum absolute error decreases to zero very quickly. This demonstrates the usefulness of the optimality of these filters with respect to the sup-norm and $l_{2}$-norm compared to other polynomial filters.
VII-B Glauber Chain
In this section, we consider a Glauber chain on a cycle [21]. Let $p$ be the number of vertices in the cycle. Each vertex in this graph can take two values: $+1$ or $-1$. Hence, the state space of this Markov chain is ${\mathsf{X}}\coloneqq\{+1,-1\}^{p}$. In Figure 3, an instance of a state transition of this Markov chain is given where $p=11$. In these two cycles, only the direction of the top nodes are different, and so, with positive probability, a transition can occur.
$\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$
$\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$
Figure 3: A state transition of a Glauber chain on the cycle. Up arrow means $+1$, down arrow means $-1$.
In this model, the Gibbs distribution is given by
$\pi(x)=\frac{e^{-\beta H(x)}}{Z(\beta)},$
where $H(x)=-\sum_{u,w\in\mathbb{Z}_{p}:u\sim w}J_{v,w}\,x(v)x(w)$ is the energy of the state $x$ and $Z(\beta)\coloneqq\sum_{x\in{\mathsf{X}}}e^{-\beta H(x)}$ is the partition function. Therefore, we have
[…]
where $S(x,w)\coloneqq\sum_{u:u\sim w}J_{u,w}\,x(u)$. For this Markov chain, we can obtain its second largest eigenvalue exactly via computing eigenvalues of some low-dimensional matrix [21, Corollary 6]. Indeed, let $M$ be $p\times p$ matrix whose entries are all zero except immediately above and below the diagonal and non-zero entries have the following description:
$M(i,i-1)$ $=s_{i-1}/(c_{i-1}+c_{i})$
$M(i,i+1)$ $=s_{i}/(c_{i-1}+c_{i}),$
where $J_{i}\coloneqq J_{i,i+1}$, $s_{i}\coloneqq\sinh(2J_{i})$, and $c_{i}\coloneqq\cosh(2J_{i})$. Let the eigenvalues of $M$ be $\gamma_{1}\geq\gamma_{2}\geq\ldots\geq\gamma_{p}$. Then, by [21, Corollary 6], the second largest eigenvalue of $P$ is given by
$1-\frac{1-\gamma_{1}}{p}.$
Hence, we can take $\lambda_{{\mathsf{low}}}$ as
$\lambda_{x_{**}}=\frac{1-\gamma_{1}}{p}\eqqcolon\lambda_{{\mathsf{low}}}.$
In the numerical experiment, we take $p=4$, $J_{i}=1$ for all $i$, and $\beta=0.2$. We generate a random function $f$ as follows:
$f=[9.04,9.79,4.38,1.11,2.58,4.08,5.94,2.62,6.02,7.11,$
2.21,1.17,2.96,3.18,4.24,5.07 2.21,1.17,2.96,3.18,4.24,5.07]2.21 , 1.17 , 2.96 , 3.18 , 4.24 , 5.07 ]
In this case, $\lambda_{{\mathsf{low}}}=0.155$. We apply the ergodic average, Bernstein polynomial filter, Chebyshev polynomial filter, and Legendre polynomial filter up to degree $20$, and obtain the results depicted in Figure 4.
Figure 4: A maximum absolute error as a function of degree of the filter.
In this figure, the maximum absolute error is computed as in the previous example. It is evident that the Bernstein polynomial filter exhibits a slight improvement over the ergodic average. However, both the Chebyshev and Legendre polynomial filters markedly surpass the ergodic average, as anticipated. In fact, for these polynomials, the maximum absolute error diminishes rapidly towards zero similar to the previous example.
VIII Conclusion
In this paper, we explore the optimization of ergodic or Birkhoff averages in the context of the ergodic theorem for reversible Markov chains via utilizing graph signal processing techniques to achieve faster convergence to desired values. We use the transition probabilities of a Markov chain to form a directed graph so that we can represent a function on the state space as a graph signal. This allowed us to introduce a graph variation concept and define the graph Fourier transform for graph signals on this directed graph. We identify the inherent non-optimality of the standard graph filter used in ergodic theorem iterations and address this by formulating three optimization problems with distinct objectives. The solutions to these problems yield the Bernstein, Chebyshev, and Legendre polynomial filters, each offering improved performance over traditional method. Our numerical experiments demonstrate that the Bernstein filter provides a modest improvement over the ergodic average. More significantly, the Chebyshev and Legendre filters show substantial enhancements, achieving rapid convergence to the desired value. These results highlight the potential of using optimized graph filters to improve the efficiency of ergodic averages.
The success of the proposed filters in this study shows the power of graph signal processing in addressing complex problems in stochastic processes and dynamical systems. Future work could extend this approach to other types of Markov chains and more abstract state spaces. Additionally, investigating the practical applications of these optimized filters in various domains, such as statistical physics, algorithm design, and network theory, could provide valuable insights and benefits.
IX Acknowledgements
The author is grateful to Professor Alexander Goncharov for directing him to significant results in approximation theory.